# Learning in a Multifield Coherent Ising Machine

### v1

@@ Abstract
Physical information processors can learn from examples if they are modified according to an abstract parameter update equation, termed a learning rule. We introduce a physical model for self-learning that encodes the learning rule in the Hamiltonian of the system. The model consists of a network of multi-modal resonators. One of the modes is driven parametrically into a bi-stable regime, forming a coherent Ising machine (CIM)—that provides the long-term memory that stores learned responses (weights). The CIM is augmented with an additional spinor field that acts as short-term (activation) memory. We numerically demonstrate that, in the presence of suitable nonlinear interactions between the long-term memory Ising machine and the short-term memory auxiliary field, the system autonomously learns from examples.
Physical computing, parametric oscillators, physical learning, tight-binding model
preprint: APS/123-QED
@@ Introduction
Physical computing [1]—computing based on novel and unconventional phenomena and information carriers, has achieved remarkable successes. Some examples are photonic image recognition [2, 3] and equation solving [4, 5]; the use of thermal noise for matrix inversion [6], to perform inference [7] and to sample probability distributions [8]; the application of condensed-matter ideas such as topology [9], symmetry [10] and parametric phenomena [11] for computing; and the use of mechanical information carriers for the realization of acoustic arithmetic operations [12], passive speech recognition [13], passive event counting [14], and to embody intelligent responses in soft robots [15]. In contrast, physical learning—the process by which a physical computer autonomously improves its performance at a task by updating its parameters according to feedback and examples—is still comparatively in its infancy. Remarkable steps have been taken with the discovery of learning rules [16, 17, 18, 19, 20, 21, 22, 23]—that indicate how parameters should be updated to improve system performance in a given information processing task; and some of these learning rules have been demonstrated in table-top experiments [24, 25, 26]. However, a significant gap remains between learning rules—which are abstract parameter update equations, and physical systems—governed by a Hamiltonian, and its interaction with the environment in the form of fluctuations and dissipation.
Figure 1: A multifield coherent Ising machine is a network of multi-modal resonators, represented by a tight-binding model consisting of a scalar field $\sigma$ interacting with a spinor field $\vec{\psi}$. At every site $i$, the model has three degrees of freedom $\psi_{x,i}$, $\psi_{y,i}$ and $\sigma_{i}$, coupled via the nonlinear terms $H_{L}$ and $H_{I}$. The field $\vec{\psi}$ corresponds to the computational degrees of freedom (neural activations). $\vec{\psi}$ is composed of two identical sub-systems $\psi_{x}$ and $\psi_{y}$ (mass-spring lattices). During training, the output degree of freedom of the sub-system $\psi_{y}$ is clamped to a target value ($\psi_{x}$ is left free). The modes $\sigma_{i}$ are parametrically pumped (dotted green line) into a bistable self-oscillation regime, modeled by a spin (up or down depending on the phase of oscillation). This results in an Ising-like system that provides a long-term memory and stores the learned weights. The nonlinear interaction term $H_{I}$ modulates the dynamics of the computational field $\vec{\psi}$ according to the weights $\sigma$ (resulting in strong springs between antiferromagnetically-aligned sites). The term $H_{L}$ causes the weights $\sigma$ to be updated in response to a difference (contrast) between the free ($\psi_{x}$) and clamped ($\psi_{y}$) computational sub-systems. In a material, the field $\vec{\psi}$ can be realized by mapping the two spin components to a pair of doubly-degenerate localized orbitals.
In this letter, we introduce a material model—a tight-binding Hamiltonian—that learns autonomously from a set of examples (Fig. 1). The model harnesses thermal fluctuations to drive transitions between states, and uses symmetry-protected degeneracies to construct two copies of the system, as required by contrastive learning rules. The incorporation of degeneracy and noise as building blocks of a self-learning physical system illustrates how phenomena that are not traditionally used for computation—or that are even actively avoided—can be reclaimed in novel information processing paradigms. The proposed system, referred here as a multifield coherent Ising machine (Fig. 1), consists of a network of multi-modal resonators. Every site has three modes $\sigma$, $\psi_{x}$ and $\psi_{y}$. One of the modes, termed the long-term memory, learning or spin mode $\sigma$, is parametrically pumped into a bi-stable regime, forming a synthetic spin system known as a coherent Ising machine (CIM) [27, 28]. We use the bi-stability of the spin mode $\sigma$ to store learned responses over long time scales, with the spin texture playing the role of the weights in a machine learning model. In addition to the spin field $\sigma$, two degenerate resonator modes are used to construct a synthetic spinor field $\vec{\psi}=(\psi_{x},\psi_{y})$. These modes will become the computational or short-term memory degrees of freedom—holding intermediate results of computations (neural activations). Nonlinear interactions between the modes, encoded in a learning Hamiltonian $H_{l}$, will cause the short-term dynamics of $\vec{\psi}$ to be conditioned by the long-term memory $\sigma$—allowing the field $\sigma$ to parameterize the computation. $H_{l}$ will also cause the weights in $\sigma$ to be updated according to a learning rule. A CIM interacting with an additional spinor field $\vec{\psi}$ through a nonlinear term $H_{l}$ is described by a Langevin equation:
@@
Where $i$ is the site number. We set the frequency of the computational and learning degrees of freedom to $\omega_{c}^{2}=0.5$ and $\omega_{l}^{2}=1$ respectively. The corresponding quality factors are set to $Q_{c}=4000$ and $Q_{l}=40$, while the pumping strength is set to $\alpha_{l}=0.1$ and the Duffing coefficient is set to $\epsilon=0.005$. The terms $\xi$ correspond to uncorrelated (diagonal), Gaussian distributed thermal noise with an autocorrelation $\langle\xi_{l/c}(t_{0})\xi_{l/c}(t_{1})\rangle=\sqrt{2k_{B}Tb_{c/l}}\delta(t_{1}-t_{0})$. Interactions between degrees of freedom are captured by the learning potential $H_{l}=H_{I}+H_{L}$, where the local part $H_{L}$ couples the fields locally at every site of the lattice—updating the long-term memory according to the learning rule—and the interaction part $H_{I}$ couples each site with its nearest neighbors—ensuring that information is processed according to the weights stored in the $\sigma$ field.
Figure 2: Single-site learning dynamics. (a) Parametric self-oscillation regions corresponding to the contrast $\psi_{X}-\psi_{Y}$ values of 00 (red), $3$ (purple) and $5$ (brown). These contrasts are the left limit, dashed vertical line and right limit in panel (b). The red dot corresponds to the excitation parameters (frequency of $22\omega_{l}=$ 2 and amplitude of $\omega_{l}^{2}\alpha=0.1$) (b). b Probability of flip during a learning protocol as a function of the contrast $\psi_{X}-\psi_{Y}$ for $k_{B}T$ values of $0.001$ (blue), $0.05$ (orange) and $0.1$ (green). c Transient dynamics during a learning protocol of width $\Delta_{T}=300$. The top panel shows the interaction strength $\mu(t)$ during the protocol. The middle panel shows the energy in the Ising mode $\sigma$ (denoted $E_{I}$), corresponding to a contrast of 00 (red), $3$ (purple) and $5$ (brown), with a temperature $k_{B}T=0.05$. The black curve corresponds to a contrast of $5$ but setting $K_{B}T=0.05$. The bottom panel shows the Rabi oscillations in the energy $E_{X}$ of the free ($\psi_{x}$) mode, and the energy $E_{Y}$ of the clamped ($\psi_{y}$) mode, during the learning pulse. d Probability of spin flip as a function of the contrast $\psi_{x}-\psi_{y}$, for learning potentials $\mu_{m}=\mu_{m,0}/n^{2}$ with $n=1$ (blue), $n=2$ (orange) and $n=3$ (green) [top] and of protocol durations $\Delta_{t}=50$ (blue), $\Delta_{t}=100$ (orange), and $\Delta_{t}=300$ (green) [bottom]. The dashed lines in the top panel shows the sigmoidal fit of Eq. 4. Throughout the figure, $\mu_{m,0}=0.005$.
The learning Hamiltonian $H_{l}$ is designed by analogy to the contrastive learning rule [24], which involves comparing two copies of a system. We refer to these copies as free and clamped sub-systems. Both copies are presented with an excitation (in this work, a harmonic force at the resonance frequency $\omega_{c}$ of the computational modes $\psi_{x/y}$), at one or several input sites. The amplitude of this force encodes the input values. The output is obtained by reading the amplitude of oscillation of an output site. During learning, the output site of the clamped sub-system is forced to take a specific vibration amplitude. Then, the parameters (springs) are updated proportionally to the difference between the clamped and free energies. Although it can be proven that contrastive learning approximates gradient descent, intuitively, the learning rule can be understood as the changing the model parameters until the free system behaves like the clamped one—thus learning to produce the desired output.
The need for two system copies poses a practical challenge for implementing contrastive learning rules. In our realization, we address this by mapping each copy to a component of the spinor field $\vec{\psi}$, with $\psi_{x}$ and $\psi_{y}$ becoming the free and clamped copies respectively. Although our work concerns a tight-binding model, in an actual condensed matter system, the two identical copies of the system can be robustly realized by associating them with basis functions that transform under a two-dimensional irreducible representation of the system’s symmetry group, mapping each copy $\psi_{x}$ and $\psi_{y}$ to an irreducible symmetry sector [29]. In contrastive learning, springs are updated in response to a difference between forces in the clamped and free configurations. Here, we implement a variation of the learning rule where the long-term memory is updated (spins are flipped) in response to the clamped and free site oscillation amplitudes. Although our learning rule deviates significantly from traditional contrastive learning, as spins can only take discrete states and experience probabilistic updates, a similar intuition as with contrastive learning can be applied: spins will flip until the free copy—which does not have access to the target output—learns to naturally approximate the target values. This intuition will be validated numerically in the next sections.
@@
Our learning rule boils down to a conditional spin forgetting mechanism: the phase of a spin $\sigma_{i}$ is randomized with a probability conditional on the difference between the clamped and free configurations at the site, thereby inducing spin flips in sites where free and clamped are misaligned. This mechanism arises from the following nonlinear, local interaction term acting on each site $i$,
$H_{L,i}=\mu(t)\left(\psi_{x,i}-\psi_{y,i}\right)^{2}\sigma^{2}_{i}=\mu(t)\psi_{x-y,i}^{2}\sigma_{i}^{2}.$ (3)
Since $\psi_{x}$ and $\psi_{y}$ are degenerate modes, the contrast between clamped and free configurations $\psi_{x-y}$ is merely a rotation of the local basis $(\psi_{x+y},\psi_{x-y})=(\hat{\sigma}_{x}+\hat{\sigma}_{z})(\psi_{x},\psi_{y})$, where $\hat{\sigma}_{x/y/z}$ is the corresponding Pauli matrix. This allows us to interpret $H_{L}$ as a cross-Kerr interaction between the Ising field $\sigma$ and the rotated computational field $\psi_{x-y}$. Thus, the term introduces a change in the resonance frequency of the mode $\sigma_{i}$ that increases quadratically with the contrast $\psi_{x-y}=\psi_{x}-\psi_{y}$ between clamped and free configurations. If $\psi_{x-y}$ is sufficiently large, the system will be detuned outside the parametric resonance region (Fig. 2a), causing its oscillation to decay, thereby forgetting its state (Fig. 2b,c). At low temperatures, the detuning results in a sharp, step-like transition: If $\psi_{x-y}$ is above a certain threshold, the system forgets its state, resulting in a probability of flip of $1/2$, while nothing occurs below the threshold. For higher temperatures, the probability of a spin flip becomes a smooth function of the contrast $\psi_{x-y}$ (Fig. 2b).
@@
Figure 3: Learning in a 4x4 lattice (a) System under consideration. A harmonic force with frequency $\omega_{c}$ and amplitude $F_{i}$ is applied at the top-left site, to both clamped and free degrees of freedom. The output amplitude $A_{o}$ is measured at the bottom-right corner. (b) Cumulative density of states (orange) and density of states (blue) as a function of the transmissivity, computed via linearized model. The dots correspond to simulations of the transmissivity using the full nonlinear ODE with an excitation force of $10^{-4}$. (c) Spin textures corresponding to the dots in (b) (d) Evolution of the transmissivity as a function of training iteration for target values $A_{T}$ of 00, $1$ and $2$ (dashed line). The line represents the mean of N simulations, while the shaded area represents a standard deviation. (e) Output amplitude $A_{o}$ as a function of the clamping amplitude, after 200 learning iterations, starting from a random configuration. The shaded area represents one standard deviation. The dashed orange line corresponds to an ideal learning response. We observe that, for conductivity values for which spin state exists, the transmitted amplitude approximately converges to the clamping amplitude. Panels (d) and (e) are calculated with a sigmoidal flip probability ($p_{0}=0$, $\psi_{T}=0.16$ and $\beta=1500$) and an input harmonic force $F_{i}=10^{-3}$.
An unwanted side effect of the potential in Eq. 3 is that it couples the free and clamped sub-lattices. This is exemplified by the presence of Rabi oscillations between $\psi_{x}$ and $\psi_{y}$ when the potential coefficient $\mu(t)$ is nonzero (Fig. 2c). However, the learning rule requires comparing two independent lattices. We mitigate this effect by keeping the interaction strength to zero, $\mu(t)=0$, during equilibration of the field $\vec{\psi}$ in response to a change in the input. At every learning iteration, we apply the learning potential according to a Gaussian pulse, characterized by a peak value $\mu_{m}$ and a width $\Delta_{t}$, $\mu(t)=\mu_{m}\exp{\left(-\frac{(t-t_{0})^{2}}{2\Delta_{t}^{2}}\right)}$ ). This potential causes the spins $\sigma$ to be updated in response to the difference between clamped and free configurations. Since the decay rate of the the Ising mode $\sigma_{i}$ is much faster than the decay rate associated with the computational modes $\vec{\psi}_{i}$, the contrast $\psi_{x-y}$ remains significant during the learning iteration (Fig. 2c). Even if Rabi swaps occur between $\psi_{x}$ and $\psi_{y}$, the contrast $\psi_{x-y}$ is unaffected as $H_{L}$ does not mix the $\psi_{x+y}$ and $\psi_{x-y}$ sectors.
Intuitively, we would expect the probability of spin-flip to follow a step-like change as a function of $\psi_{x-y}$: Below a critical detuning, the system remains in resonance and remembers its state, while above the threshold $\psi_{T}$ the system falls out of resonance and forgets its state (Fig. 2a). At finite temperatures, this step-like response becomes a smooth sigmoidal function (Fig. 2b,d):
$p(\psi_{x}-\psi_{y})=p_{0}+\frac{1}{4}\left[1+\text{tanh}\left(\beta(\psi_{x-y}^{2}-\psi_{T})\right)\right],$ (4)
in which the threshold $\psi_{T}$ and the slope $\beta$ are fitted empirically to a numerical simulation (Fig. 2d). Remarkably, the sigmoidal description breaks down for shorter pulse widths $\Delta_{t}$. In these cases, the oscillation of $\sigma$ does not have time to decay below the thermal noise, thus forgetting its phase state (spin). In this short-pulse regime, the system presents coherent bit flipping—the spin changes deterministically for specific values of $\psi_{x-y}$—as opposed to conditional forgetting (Fig. 2d), which is the intended response.

### v2

@@ Abstract
We introduce a network of coupled oscillators that can learn to solve a classification task from a set of examples—performing both training and inference through the nonlinear evolution of the system. We accomplish this by combining three key elements to achieve learning: A long-term memory that stores learned responses, analogous to the synapses in biological brains; a short-term memory that stores the neural activations, similar to the firing patterns of neurons; and an evolution law that updates the synapses in response to novel examples, inspired by synaptic plasticity. Achieving all three elements in wave-based information processors such as metamaterials is a significant challenge. Here, we solve it by leveraging the material multistability to implement long-term memory, and harnessing symmetries and thermal noise to realize the learning rule. Our analysis reveals that the learning mechanism, although inspired by synaptic plasticity, also shares parallelisms with bacterial evolution strategies, where mutation rates increase in the presence of noxious stimuli.
@@ Introduction
Wave-based computing offers the tantalizing prospect of massive parallelism [1], speed-of-light operation [2] and low energy consumption [3]. Applications span a broad range of problems, including evaluating arithmetic operations [4], computing derivatives [5], solving integral equations [6], performing voice recognition [7, 8], image classification [9, 10], and generating pictures through reverse diffusion processes [11, 12]. Traditionally, metamaterials that compute with waves are designed for fixed tasks via top-down algorithms such as adjoint optimization [13, 14] or combinatorial methods [15]. This top-down approach contrasts with the continuous learning process that is characteristic of living brains, in which connections between decentralized neurons are autonomously formed using local information [16, 17, 18]. This decentralized learning process confers a resilience and adaptability that artificial systems lack, and reduces training costs.
Physical learning is an emerging field that aims at bringing the adaptability of biological organisms to artificial materials, by creating synthetic systems that learn from examples. A key strategy in physical learning consists in updating the model parameters according to a learning rule. These local learning rules are inspired by the synaptic plasticity mechanisms in neurons, with Hebbian learning [16, 17]—the notion that neurons form connections based on correlations between firing patterns—being one of the foremost examples. Yet learning strategies also include many recent additions, such as equilibrium propagation [19], the forward-forward algorithm [20] and Hamiltonian echo backpropagation [21]. While much of the work has focused on developing and analyzing learning rules in abstract or algorithmic terms, an equally important frontier lies in the physical realization of systems that embody these rules. In this context, self-learning systems have been realized by externally updating the system parameters according to a learning rule, either manually [22] or using electronic circuits [23].
Figure 1: Self-learning metamaterial (a) During inference, the metamaterial acts as a mass-spring network. Features are encoded in the excitation amplitude at designated input sites (yellow and purple spheres). If the output amplitude exceeds a threshold, an object is recognized. (b) For the material to learn according to a contrastive learning rule, an identical copy to the lattice in (a) is constructed. The original lattice is operated normally, letting the output site vibrate freely, and is referred to as the free lattice. In the copy, the output is clamped to vibrate at the correct amplitude. At every learning iteration, the coupling springs are updated depending on the difference between the vibrating amplitude of the clamped and free lattices. (c) In the proposed metamaterial, the masses of the free and clamped lattices are mapped to normal modes of each metamaterial site. We refer to these as $\psi_{x}$ and $\psi_{y}$ respectively (collectivelly referred as computational degrees of freedom). The coupling springs are encoded in an additional mode, that is represented by $\sigma$. This mode is parametrically driven at twice its natural frequency (dotted green line) and consequently has two stable phases of oscillation (represented as spin states with blue and red arrows). As a consequence of a nonlinear interaction potential, the effective springs connecting the computational modes depend on the weight state—taking a weak value when neighboring strings are aligned, and a strong value when neighboring weights are oscillating in opposite phases. The learning potential (blue line connecting the modes) causes the weight state to change when clamped and free copies disagree, implementing a contrastive learning rule.
An open problem in physical learning is figuring out how to build physical systems where the learning occurs through the natural evolution of the material, instead of relying on manual updates or external electronic circuitry. Such material would extend the parallelism and speed of wave computing to the learning process, enabling the use of physical phenomena to train machine learning models. However, realizing self-learning in metamaterials is a significant challenge. This is because a self-learning material is tasked with solving two simultaneous problems [24]: On one hand, the material performs a direct problem on the input signals. This direct problem can be, for example, solving a classification task. On the other hand, the material is also performing an inverse problem, autonomously adapting its parameters to improve its fitness in response to new examples.
Results
In this work, we introduce a metamaterial model that learns from examples. The metamaterial, consisting of a network of coupled multimodal resonators, simultaneously solves both learning and inference problems using wave physics. When solving a classification task, the proposed metamaterial can be seen as a square lattice of masses and springs. The input is applied as a harmonic force in one or several sites. In a flower classification problem, this input may consist of the geometric features of the flower, such as petal width and petal length—with higher amplitudes corresponding to larger feature values. The output of the computation is the vibration amplitude of a designated output site (Fig. 1a). Although spring networks (and mathematically analogous resistor networks) are relatively simple physical systems, they are able to solve classification [25] and allostery [26] tasks; when they are augmented with nonlinearity, they are in fact computationally universal [27]. The learning process is based on the contrastive learning rule [28]. In contrastive learning, two copies of the system are constructed. Both copies are excited with the same input. During training, the output of one of the copies is clamped to the target value, while the other copy is left free. Then, the parameters (springs) are updated proportionally to the difference between vibration amplitudes in the clamped and free copies (Fig. 1b). This learning process can be intuitively understood as changing the system parameters until the free copy—which has not been exposed to the correct answer—learns to mimic the copy that is clamped to the correct result.
The requirement for two copies of the system has been an obstacle to the realization of contrastive learning schemes [29]. Here, we overcome this challenge by relying on the crystalline symmetries of the metamaterial (Fig. 1c). When a lattice exhibits a symmetry, such as the four-fold rotational symmetry $C_{4}$ present in our case, eigenmodes corresponding to two-dimensional irreducible representations of the symmetry automatically appear in pairs [30]. Around the frequency of these eigenmodes, the system effectively behaves as two separate, identical oscillator networks. These two networks can be addressed independently by controlling the location at which excitation and clamping are applied. We will use these two symmetry-protected, identical copies to construct the free and clamped copies required by the contrastive learning rule. We will denote the modal displacements at site $i$ as by ${\psi_{i}}_{x}$ and ${\psi_{i}}_{y}$ respectively, and refer to both of them together as the computational field $\vec{\psi}$.
In self-learning mass-spring networks, the values of the springs connecting the computational degrees of freedom are also dynamical variables. To allow for the spring constants to vary, we encode them in an additional site eigenmode, that we refer to as the weight field $\sigma_{i}$. While the computational field $\vec{\psi}$ is different for every input sample—for example, for every flower being classified—the weight field $\sigma_{i}$ accumulates all prior learned experience of the metamaterial. Thus, it requires a much longer memory time scale. We realize this long-term memory by parametrically driving the mode, modulating its local stiffness at two times the modal resonance frequency. When a system is parametrically driven strongly enough, the system undergoes a bifurcation from which two stable, self-oscillating solutions emerge. These two stable solutions—characterized by opposite phases of oscillation—can be extremely long-lived, with transitions driven by thermal noise. Since the two stable states can be modeled by a spin-like binary variable, the weight field $\sigma$ can be thought of as emulating an Ising model. Such emulators are commonly referred to as coherent Ising machines.
The self-learning metamaterial can be seen as a coherent Ising machine $\sigma_{i}$ interacting with the computational field $\vec{\psi}_{i}$. Thus, we refer to it as a multifield coherent Ising machine. Throughout the rest of this section we will discuss how the learning rule emerges from nonlinear interactions between computational and weight fields. The nonlinearity plays a double role. First, it makes the springs that connect computational degrees of freedom dependent on the weight field $\sigma$. Second, it modulates the long-term memory of the weight field $\sigma$ depending on the difference between clamped and free computational fields—causing the system to ‘forget’ the weights on sites where clamped and free copies disagree. We will finally show how these two interactions together allow the system to learn a classification problem.
The equations of motion of the material, describing the evolution of the fields $\vec{\psi}$ and $\sigma$ at each site $i$ are:
@@ Results
We set the frequency of the computational and learning degrees of freedom to $\omega_{c}^{2}=0.5$ and $\omega_{l}^{2}=1$ respectively. The corresponding quality factors are set to $Q_{c}=4000$ and $Q_{l}=40$, while the parametric driving strength—how much the resonator frequency is externally varied over time—is set to $\alpha_{l}=0.1$. The nonlinear Duffing coefficient plays a critical role in preventing the amplitude of the mode $\sigma_{i}$ to grow without bounds, and is set to $\epsilon=0.005$. The terms $\xi$ represent the thermal noise exerted by the environment on the system. They take the usual form of a set of independent, Gaussian distributed random forces with autocorrelation $\langle\xi_{l/c}(t_{0})\xi_{l/c}(t_{1})\rangle=\sqrt{2k_{B}Tb_{c/l}}\delta(t_{1}-t_{0})$. Interactions between degrees of freedom are captured by the learning potential $H_{l}=H_{I}+H_{L}$, where the local part $H_{L}$ couples the fields locally at every site of the lattice—updating the long-term memory according to the learning rule—and the interaction part $H_{I}$ couples each site with its nearest neighbors—ensuring that the field $\vec{\psi}$ processes information according to the weights stored in the $\sigma$ field.
Figure 2: Single-site learning dynamics. (a) Evolution of the energies in the weight ($\sigma$, middle panel) and computational ($\psi_{x}$, $\psi_{y}$, bottom panel) resonators, as a learning pulse is applied (top panel). b Probability of flip during a learning protocol as a function of $\psi_{X}-\psi_{Y}$ for different temperatures, corresponding to $k_{B}T$ values of $0.001$ (blue), $0.05$ (orange) and $0.1$ (green). c Parametric self-oscillation region (shaded grey area) of the weight degree of freedom $\sigma$; in the shaded region, the weight is bistable and thus has long term memory. The excitation conditions are shown as dashed orange lines and red dot. When $\psi_{X}-\psi_{Y}$ is not zero, applying the learning potential shifts the self-oscillation region. The value of $\psi_{x}-\psi_{y}$ where $\sigma$ goes out of parametric resonance is shown as a purple dashed line in panel b. d Probability of spin flip as a function of $\psi_{x}-\psi_{y}$, for learning potentials $\mu_{m}=\mu_{m,0}/n^{2}$ with $n=1$ (blue), $n=2$ (orange) and $n=3$ (green), illustrating how the step location can be shifted by setting the potential. e Probability of spin flip as a function of the protocol durations $\Delta_{t}=50$ (blue), $\Delta_{t}=100$ (orange), and $\Delta_{t}=300$ (green), showing the emergence of probability oscillations at short learning pulses. The dashed curves in panels c and d show the fit with Eq. 5. Throughout the figure, the parameters are $2\omega_{l}=2$, $\omega_{l}^{2}\alpha=0.1$, $\mu_{m,0}=0.005$, and the step duration is $\Delta_{T}=300$ unless stated otherwise.
@@
In the proposed metamaterial, the learning rule arises due to local interactions at each site $i$. These interactions update the weight field ($\sigma$) when the free ($\psi_{x}$) and clamped ($\psi_{y}$) sub-systems disagree. In traditional contrastive learning, the weights are continuous variables, adjusted in proportion to the difference between clamped and free configurations $\psi_{x}-\psi_{y}$. However, in our model, every site of the weight field has only two stable phases of oscillation and, therefore, cannot be varied continuously. We adapt to this difference by implementing a probabilistic update process, where the probability that a weight $\sigma_{i}$ flips increases with increasing $|\psi_{x}-\psi_{y}|$. To realize this probabilistic bit flipping, we rely on the fact that the memory of a parametric oscillator—the average time it takes for the state to flip randomly due to thermal noise—is highly dependent on the detuning [31]—the difference between its parametric excitation frequency and the optimal parametric resonance condition. We can thus realize probabilistic contrastive learning by applying a potential that shifts the natural frequency of $\sigma_{i}$ as a function of $\psi_{y}-\psi_{x}$, causing the system to be detuned.
Such selective frequency shift can be accomplished with a cross-Kerr interaction. Cross-Kerr interactions shift the frequency of one mode as a function of the energy in another mode, and take the form
$H_{L,i}=\mu(t)\psi_{x-y,i}^{2}\sigma_{i}^{2}=\mu(t)\left(\psi_{x,i}-\psi_{y,i}\right)^{2}\sigma^{2}_{i}.$ (3)
Here, to turn a cross-Kerr interaction into a difference-sensitive detuning, we used the rotational symmetry of the site. Since the sites are symmetric, any combination between the eigenmodes $\psi_{x}$ and $\psi_{y}$ is also an eigenmode. Thus, we can look at the rotated set of modes $(\psi_{x+y},\psi_{x-y})=(\psi_{x}+\psi_{y},\psi_{x}-\psi_{y})$. In this new basis, the difference between modes $\psi_{x}-\psi_{y}$ can be seen as its own mode $\psi_{x-y}$. In an experimental setting, such a rotated cross-Kerr interaction could be realized by introducing a non-linear spring in a region of the resonator where the two modes $\psi_{x}$ and $\psi_{y}$ have opposite signs.
The potential in Eq. 3 breaks the $C_{4}$ rotational symmetry of the resonator; when this term is active, the system can no longer be seen as two separate copies. We overcome this by setting the strength $\mu(t)$ to zero during computation, preserving the two-copy picture. To initiate a weight update, we briefly turn on the cross-Kerr term, by increasing $\mu(t)$ according to a Gaussian pulse with peak value $\mu_{m}$ and a width $\Delta_{t}$,
[…]
Figure 2a shows the evolution of the energies in the modes $\sigma$, $\psi_{x}$ and $\psi_{y}$ during a weight update. The resonator is initialized with specific values of the fields $\vec{\psi}$ and $\sigma$. During operation, the initial condition for $\vec{\psi}$ is set by the interaction with the neighbors, as will be discussed later in the paper. Our simulations reveal that, when the learning potential is applied, the oscillation of the weight degree of freedom ($\sigma)$ decreases with increasing difference between initial conditions of $\psi_{x}$, $\psi_{y}$. This is a consequence of the increased detuning of $\sigma$, which drives the mode away from resonance. The lower amplitude of the detuned mode makes thermally induced flips more likely, thus reducing the memory of the system. When the learning potential is active, we also observe a periodic exchange of energy between free and clamped modes. This is an expected consequence of $H_{L}$ breaking the fourfold rotational symmetry—introducing a difference between the $x+y$ and xyx-y directions and thus distorting the picture where the clamped and free copies are independent.
The difference between clamped and free oscillations affects the probability of flipping a weight in a step-like fashion (Fig. 2b). This step-like response arises because $\sigma_{i}$ only self-oscillates in a specific region of parametric excitation frequencies and amplitudes (Fig. 2c). When the difference between clamped and free oscillators is large enough, the nonlinear interaction (Eq. 3) shifts this region outside of the excitation conditions. As a consequence, the system ceases to self-oscillate and forgets its state. At low temperature, this phenomenon is abrupt—as long as the excitation remains within the self-oscillation region, the system remembers its state. In contrast, when the temperature is finite, random thermal flips start to occur at finite amplitudes, smoothing out the response (Fig. 2b). The difference $\psi_{x}-\psi_{y}$ at which the forgetting transition occurs is a tunable parameter. We can control it by adjusting the strength of the interaction $\mu_{m}$ (Fig. 2d).
@@
Figure 3: Site-site interactions in a lattice (a) Example metamaterial, consisting of a 4x4 lattice. A harmonic force with frequency $\omega_{c}$ and amplitude $F_{i}$ is applied to both clamped and free degrees of freedom at the input site, situated at the top-left corner. The output amplitude $A_{o}$ is measured at the bottom-right corner. (b) Cumulative density of states (orange) and density of states (blue) as a function of the transmissivity, computed via the effective mass-spring model. The dots correspond to simulations of the transmissivity using the full nonlinear system (Eq. 2) with an excitation force of $10^{-4}$. (c) Weight configurations corresponding to the dots in (b) (d) Evolution of the transmissivity as a function of training iteration for target values $A_{T}$ of 0, 11 and 22 (dashed line). The line represents the mean of N simulations, while the shaded area represents a standard deviation. (e) Output amplitude $A_{o}$ as a function of the clamping amplitude, after 200 learning iterations, starting from a random configuration. The shaded area represents one standard deviation. The dashed orange line corresponds to an ideal learning response. We observe that, for transmisivitty values for which a weight configuration exists, the transmitted amplitude approximately converges to the clamping amplitude.
We empirically observe that the probability of transitions $p$ closely follows (Figs. 2b,d) a hyperbolic tangent relation,
$p=p_{0}+\frac{1-2p_{0}}{4}\left[1+\text{tanh}\left(\beta(|\psi_{x-y}|-\psi_{T})\right)\right],$ (5)
in which $\psi_{T}$ is the threshold for bit flipping, $\beta$ determines the slope of the flipping transition, and $p_{0}$ is the probability of spontaneous thermal flipping in absence of contrast between clamped and free configurations. This is unsurprising, because our weight field behaves as an effective spin system, and hyperbolic tangents commonly arise in spin-flip probabilities due to detailed balance considerations, as exemplified paradigmatically by Glauber dynamics [32]. Later, Eq. 5 will allow us to predict the probability of bit flips in large lattices without having to simulate expensive stochastic differential equations for each site. Remarkably, the sigmoidal description breaks down for shorter pulse widths $\Delta_{t}$. In these cases, the oscillation of $\sigma$ does not have time to decay below the thermal noise, thus forgetting its state. The remaining amplitude is phase-shifted by the learning potential, giving rise to oscillations of the bit-flip probability as a function of $\psi_{x}-\psi_{y}$ (Fig. 2e) when the accumulated shift is a multiple of $\pi$. In the rest of this work, we will operate in the sigmoidal regime—corresponding to long learning pulses—to achieve a monotonic relation between weight flip probability and clamped-free difference.